How can I get the answer underlined in this case by using boolean algebra simplification?
Boean aLteba smpen
Answer #1

First consider given boolean algebra equation for doing simplification

F=AB-C-+A-B-C+A-D- = D-+A+B-C   

Take left side equation for getting right side simplified form

L.H.S=F=AB-C-+A-B-C+A-D-

=AB-C-+A-(D-+B-C)

=A+A-(D-+B-C)

=A+D-+B-C. FROM THE THEOREM A+A-B=A+B.